A plane leaves the airport on a bearing of 45degree travelling at 400 meh. The wind is blowing at a bearing of 135 degree at a speed of 40mph.what is the actual velocity and direction of the plane?
1
Expert's answer
2019-04-09T11:28:26-0400
"\overrightarrow{v}=(400\cos{45^\circ}, 400\sin{45^\circ})"

"\overrightarrow{u}=(40\cos(-45^\circ), 40\sin(-45^\circ))"

"\overrightarrow{v}+\overrightarrow{u}=(400\cos{45^\circ}+40\cos(-45^\circ), 400\sin{45^\circ}+40\sin(-45^\circ))"


"\overrightarrow{v}+\overrightarrow{u}=(440{\sqrt{2} \over2}, 360{\sqrt{2} \over2})"

"|\overrightarrow{v}+\overrightarrow{u}|=\sqrt{(440{\sqrt{2} \over2})^2+(360{\sqrt{2} \over2})^2}=40\sqrt{101} \ mph\approx402 \ mph"

"\ bearing=\arctan{440{\sqrt{2} \over2} \over 360{\sqrt{2} \over2}}=\arctan{11 \over 9}\approx50.71^\circ"